# Moving Out: Physically-grounded Human-AI Collaboration

### v2

@@ 4.3 Dataset
The data collection was approved by the Institutional Review Board (IRB). Two human players control the agents with joysticks. The game ran at 10Hz, and on average, each map took around 30 seconds (or 300 time steps) to complete the transportation of all items. See Appx. H for details.
@@
For Task 2, we emphasize the randomized properties of objects rather than the variable behaviors. In this case, we used 4 expert players to collect 720 human-human demonstrations (1,440 action sequences in total), with 60 demonstrations per map. Each map included randomized physical attributes, allowing us to evaluate the model’s ability to generalize to unseen object attributes.
@@ 5 BASS: Behavior Augmentation, Simulation, and Selection
To address the proposed tasks, we develop BASS (Behavior Augmentation, Simulation, and Selection) which considers the increased number of configurations in continuous space and the outcome of actions in physical environments. First, at training time, we augment the behavior data. This helps the model adapt to different behaviors better by exposing it to a broader range of possible interactions. Second, we train a dynamics model to simulate the outcome of an action. At inference time, the model can select actions by evaluating the predicted states. This enables the model to handle uncertainty from physical constraints and adapt its decisions accordingly.
@@ 5.1 Collaboration Behavior Augmentation
Our augmentation strategy involves two techniques:
By aligning the start and end of agent $i$’s sub-trajectory, the generated trajectories maintain temporal consistency for agent $i$ while introducing a different partner sequence. This approach enriches the training set with new, valid trajectories where agent $i$’s behavior is fixed and the partner’s varies.
Validity of the Generated States We can validate generated sub-trajectories based on the state information. Specifically, one can check whether the generated states remain within the valid state space and do not result in conflicts, such as collisions or other inconsistencies. However, we will show in the experiment section that, even without explicit validation, this augmentation strategy can improve performance.
@@ 5.2 Simulation and Action Selection
Action Selection  Once the next state is predicted, the reward for each action is computed based on the total distance of all objects to the goal region. We use Normalized Final Distance (NFD) as defined in Sec. 6, but other metrics that measure partial progress of map completion also suffice. We then select the action with the highest reward as the optimal action: $a^{*}=\arg\max_{a_{i}}r(a_{i}),i=1,2,\dots,n,$ where $r(a_{i})$ is the reward for action $a_{i}$. This approach enables the model to choose the most effective action, even in real-world scenarios without access to a simulator.
@@ 6 Experiment
We aim to answer the following research questions: (RQ1) How well do existing methods and BASS perform in physically grounded settings? (RQ2) Does BASS better support humans and work more effectively with them in physically grounded collaboration? (RQ3) What are the limitations of existing methods and BASS in physically grounded human-AI collaboration?
To answer these questions, we train and test all methods on the two Moving Out tasks to compare their performance. First, each model plays with itself. This demonstrates each model’s ability to handle physical constraints in a controlled setting (i.e., without the variability introduced by actual humans). Then, we let each model play with humans to evaluate their interactions with diverse behaviors.
@@ 6.1 Settings
  • Diffusion Policy (DP) Chi et al. (2023) captures multimodal distribution and has demonstrated strong performance across various tasks.
  • MAPPO Yu et al. (2022) is a commonly used multi-agent RL algorithm. It has demonstrated strong performance in cooperative games. See Appx. C for details about training and reward settings.
Evaluation Metrics  We measure the success of collaboration using the following metrics: (1) Task Completion Rate (TCR) for size-weighted successful item delivery; (2) Normalized Final Distance (NFD) for the distances between objects and the target, measuring partial progress; (3) Waiting Time (WT) for the amount of time an agent waits for assistance with large items; and 4) Action Consistency (AC) for the degree of force alignment when moving items jointly, indicating coordination efficiency. Detailed definitions are in Appx. F.
Human Subject Study  Our study was approved by the IRB. We conducted a human subject study with 32 participants to evaluate BASS against the DP baseline in both tasks. Each participant played 32 maps in total, cooperating with each method in two rounds per map. After completing the first round, the participant and model switched to control the other agent. Upon finishing all maps, participants were given a questionnaire to capture subjective feedback. See Appx. J for details.
Figure 6: BASS outperforms baselines on Task 1 and Task 2, against AI (solid) and humans (striped).
@@ 6.2 Results
Collaboration with AI Itself (RQ1)  Fig. 6 compares BASS with baselines in Task 1 and Task 2 by running 20 times on each map with different random seeds. BASS outperforms baselines, especially in task completion metrics TCR and NFD. For WT, DP performs better than GRU, MLP, and MAPPO, implying that it more promptly assists the other agent whenever help is needed. In terms of action consistency, there is no significant gap among methods; a potential reason is that each model is paired with itself and thus pursues the same objectives. Overall, the RL agents exhibit behaviors that follow the dense reward setting, while the BC agents mainly reflect behaviors present in the dataset. A detailed comparison of the behaviors between BC and RL agents can be found in the Appx. E.
In task 1, DP outperforms baselines in TCR and NDF, suggesting that DP effectively learns cooperative behaviors. MLP shows the longest waiting times, implying that it often leaves one agent waiting for help. By contrast, BASS reduces waiting time, likely because the scoring mechanism prioritizes overall task progress. In task 2, BASS shows stronger performance, likely because it evaluates the impact of each action on the generated state, leading to a better grasp of how to help when dealing with objects that exhibit different physical properties.
Collaboration with Humans (RQ2)  The bars with stripes in Fig. 6 show the results with humans. In both tasks, BASS significantly improved task completion rates (TCR and NFD) compared to the DP. This indicates that BASS adapts better to human behavior. For wait time, DP increased when playing with humans, suggesting it struggles with different humans, despite that DP captures multimodal distributions. BASS reduced wait time, demonstrating its ability to adapt to diverse human behaviors. For action consistency, DP performed worse because it cannot handle differences between the evaluation and training data. BASS used diverse collaborative behaviors during training and selected the best actions for interacting with humans, resulting in better consistency.
Human Feedback (RQ2)  Fig. 7 summarizes post-experiment survey results from humans. We compare BASS with DP. The results show that BASS significantly outperformed DP in the Helpfulness category, indicating that BASS is better at consciously assisting others. Additionally, BASS demonstrated a better understanding of physics, suggesting that our next state predictor effectively comprehends and evaluates different actions to choose the best ones. Independent t-tests revealed that these differences are statistically significant ($p=0.017$).
@@
Task 1 Task 2 Methods TCR↑ NFD↑ TCR↑ NFD↑ GRU 0.3070 0.3674 0.2582 0.3935 + BASS w/o Simulation 0.4117 0.4396 0.3333 0.4141 + BASS w/o Augmentation 0.3531 0.4047 0.3670 0.4246 + Full BASS 0.4120 0.4454 0.3414 0.4410 Diffusion Policy (DP) 0.3829 0.4818 0.3125 0.4526 + BASS w/o Simulation 0.4028 0.5114 0.3569 0.4908 + BASS w/o Augmentation 0.4741 0.5561 0.4200 0.5187 + Full BASS 0.5027 0.5707 0.4348 0.5535
Figure 8: Failure case study: 1) Failing to release items during handover, 2) Not responding when assistance is needed, and 3) Inability to grasp large items upon approach.
Failure case study (RQ3)  Fig. 8 shows examples of common failure cases from DP. In task 1, as illustrated in failure case 1, many participants reported that the AI agent frequently holds an item without passing it, resulting in frequent collisions. Additionally, participants noted that the AI agent often failed to come to assist, as shown in failure case 2, where a human agent (blue) was slowly pulling an item, but the AI agent (pink) instead went to grasp other smaller objects. These issues show the model’s limited ability to adapt to diverse behaviors, making it difficult to respond appropriately to actions that were not present in the training dataset. In task 2, most participants pointed out failure case 3, where the AI agent reached the target item but was unable to successfully grasp it. This indicates that the model struggles when encountering objects that were not in the training data. In contrast, BASS shows fewer reported failure cases than DP. Manual inspection revealed that BASS reduced the occurrence rates of the three failure types from {0.797, 0.688, 0.906} in DP to {0.343, 0.563, 0.484}. However, effectively addressing these failures remains a substantial challenge for future research.
@@ 7 Conclusion
We introduce Moving Out, a physically grounded human-AI collaboration benchmark that features a continuous state-action space and dynamic object interactions in a 2D physical environment. We created two challenging tasks and collected human-human collaboration data to enable future model development. Our evaluation results show that much remains to be done with existing models to effectively collaborate with humans in physical environments. Our proposed method, BASS, takes the first step to improve models’ adaptability to diverse human behaviors and physical constraints.
Limitations and Future Work  While we conduct user studies using models, achieving smooth human-AI collaboration with generative models (e.g., DP, LLMs) remains challenging because of low inference speed. In our study, it requires at least 5~10Hz to have a smooth interaction (see Appx. I.2.) Moving Out aims to include diverse physical variations and collaboration modes; however, we do not cover all possible physical interactions. Future work includes improving generative models’ inference speed to achieve smooth human-AI collaboration in physical environments, leveraging LLMs’ reasoning abilities in physically grounded collaboration tasks, and extending to more complex cooperation dynamics among multiple AI agents and humans.
@@ Appendix A Comparison with Other Environments
Environment State/Action
Physics
-based
Constraints Collaboration Behaviors
Overcooked-AI Discrete No
Items placed only
in specific locations
Passing items, dividing tasks,
and collision avoidance
Table Carrying Continuous No
No physical feedback,
task ends upon collision
Joint carrying
(i.e., action consistency)
Moving Out Continuous Yes
Realistic physics, friction,
collision feedback,
diverse items with
physical properties.
Coordination, Awareness
of needing help,
joint carrying
(i.e., action consistency).
Metrics Pros Cons Human Data
Number of cooked
onions in a limited time
Small state/action space,
fast training,
human data available
Limited behavior variety,
simple tasks
Yes
Success rate,
Completion time
Continuous actions
No physics in interactions,
single task, no dataset
No
Task Completion Rate,
Normalized Final Distance,
Waiting Time,
Action Consistency
Realistic physics,
multiple collaboration modes,
physics feedback,
human dataset available
Higher computational cost Yes
@@
Appendix B Comparison with Oracle Simulation
@@
Appendix C MAPPO Training Setting
To train MAPPO, we integrate the Moving Out environment into the BenchMARL Bettini et al. [2024] multi-agent RL library. Our approach to MAPPO training was designed to align with the objectives of Task 1 and Task 2, which were initially conceptualized with dataset-driven methods in mind. We adapted the conditions for MAPPO as follows:

### v3

@@ 4.3 Dataset
The data collection was approved by the Institutional Review Board (IRB). Two human players control the agents with joysticks. The game ran at 10Hz, and on average, each map took around 30 seconds (or 300 time steps) to transport all items. See Appx. Q for details.
@@
Dataset DTW Mean ($\uparrow$) DTW Var ($\uparrow$) Avg. Entropy (KDE) ($\uparrow$) Coverage Distance (RBF) ($\uparrow$) Moving Out Task 1 7.013 6.065 0.888 0.899 Expert dataset 4.642 3.029 0.757 0.744 RL agent collected data 4.358 2.499 0.683 0.626
Table 1: Dataset diversity across different data collection methods. Our dataset achieves consistently higher diversity compared to expert and RL agent datasets.
As shown in Table 1, we compare the diversity of our dataset against datasets collected by RL agents or experts using Dynamic Time Warping (DTW; mean and variance), entropy, and coverage distance, showing ours has best diversity. This demonstrates the effectiveness of recruiting diverse participants for data collection. See Appx. Bfor further details, including trajectory visualizations.
Figure 5: Randomization examples in Task 2, illustrating generalization to unseen physical properties.
For Task 2, we emphasize the randomized properties of objects rather than the variable behaviors. In this case, we used 4 expert players to collect 720 human-human demonstrations (1,440 action sequences in total), with 60 demonstrations per map. Each map included randomized object physical attributes, where pose, mass, and size were varied by up to 10%, while object types and shapes were randomized to be different from those used in evaluation. This setup allows us to assess the model’s ability to generalize to unseen object attributes. Fig. 5 shows examples of two maps.
@@ 5 BASS: Behavior Augmentation, Simulation, and Selection
To address the proposed tasks, we develop BASS (Behavior Augmentation, Simulation, and Selection) which considers the increased number of configurations in continuous space and the outcome of actions in physical environments. First, at training time, we augment the behavior data. This helps the model adapt to diverse behaviors better by exposing it to a broader range of possible interactions. Second, we train a dynamics model to simulate the outcome of an action, allowing the agent to understand the impact of actions on different physical properties. At inference time, the model select actions by evaluating the predicted states.
@@ 5.1 Collaboration Behavior Augmentation
Behavior Augmentation  Trajectory augmentation is already used in single-agent settings [49, 50], but extending this idea to multi-agent raises new challenges: naively altering one agent’s behavior can easily break the consistency required for cooperation, since both agents must pursue aligned goals for the trajectory to remain valid. We adopt two strategies.
First, we perturb partner poses with small Gaussian noise, $\tilde{p}_{\text{partner}}=p_{\text{partner}}+\epsilon,\quad\epsilon\sim\mathcal{N}(0,\sigma^{2}),$ while keeping other state variables unchanged, where $p_{\text{partner}}$ is the original partner’s pose, $\epsilon$ is Gaussian noise with mean 0 and variance $\sigma^{2}$, and $\tilde{p}_{\text{partner}}$ is the perturbed pose used to generate new state variations. This generates new states that mimic natural variations in human movement and improve robustness to small deviations.
Second, we augment the data by recombining sub-trajectories: the key idea is to keep one agent’s behavior fixed while swapping the partner’s subsequence, and to align the boundaries so that both agents remain goal-consistent. Formally, given a trajectory $\tau=\{(s_{t},a_{t})\}_{t=1:T}$, we extract agent-specific sequences $\tau^{i},\tau^{j}$. For time steps $t_{1},t_{2}$, we identify another trajectory $\hat{\tau}$ whose agent-$i$ states match at the boundaries, and replace agent $j$’s subsequence: $\tau^{i}\cup\bigl(\tau^{j}_{1:t_{1}-1}\cup\hat{\tau}^{j}_{t_{1}:t_{2}}\cup\tau^{j}_{t_{2}+1:T}\bigr)\cup\tau^{e}.$ This preserves agent $i$’s intention while introducing diverse yet valid partner behaviors. Further implementation details are provided in Appx. O.
Together, these augmentations expand the dataset with physically plausible trajectories that preserve collaboration-level coherence, a novelty compared to prior single-agent augmentation methods.
Validity of the Augmented States We can validate generated sub-trajectories based on the state information. Specifically, one can check whether the generated states remain within the valid state space and do not result in conflicts, such as collisions or other inconsistencies. However, as detailed in Appx. H, our recombination strategy is explicitly designed to ensure coherence (e.g., by swapping only sub-trajectories with identical start and end states). Even without additional validation, this augmentation strategy can improve performance, as we show in the experiments.
@@ 5.2 Simulation and Action Selection
Action Selection  Our policy and partner action predictor both require strong multi-modal modeling capacity to generate diverse action candidates, which forms the basis for action selection. Once the next state is predicted, the reward for each action is computed based on the total distance of all objects to the goal region. We use Normalized Final Distance (NFD) as defined in Sec. 6, but other metrics that measure partial progress of map completion also suffice. We then select the action with the highest reward as the optimal action: $a^{*}=\arg\max_{a_{i}}r(a_{i}),i=1,2,\dots,n,$ where $r(a_{i})$ is the reward for action $a_{i}$. This approach enables the model to choose the most effective action, even in real-world scenarios without access to a simulator. A comparison of NFD against alternative objectives is provided in Appx. N.
@@ 6 Experiment
We aim to answer the following research questions: (RQ1) Does BASS adapt to unseen human behaviors with limited performance degradation? (RQ2) Does BASS generalize to unseen physical constraints? (RQ3) Does BASS work more effectively with humans in physically grounded collaboration? (RQ4) What failure patterns do existing methods and BASS exhibit?
To answer these questions, we train and test all methods on the two Moving Out tasks following the designed evaluation protocols, and then conduct a human study for further validation. For AI–AI collaboration, all results are averaged over 20 runs.
@@ 6.1 Settings
  • Diffusion Policy (DP) [17] captures multimodal distribution and has demonstrated strong performance across various tasks.
  • MAPPO [30] is a commonly used multi-agent RL algorithm. It has demonstrated strong performance in cooperative games. See Appx. E for details about training.
BASS Implementation BASS builds on the same diffusion policy backbone used in our baselines, serving as both the base policy and the partner action predictor because of its strong multi-modal modeling capacity. The VAE and dynamics models are implemented as MLPs and co-trained. For action selection, the policy and partner predictor independently sample 4 action candidates each. While increasing the number of samples could further improve accuracy, collaboration requires real-time inference; sampling four candidates ensures inference can be performed at 10Hz. Ablation studies on the sampling strategy, analyses of individual modules are provided in Appx. KJ.3,  L.
Evaluation Metrics  We measure the success of collaboration using the following metrics: (1) Task Completion Rate (TCR) for successful item delivery; (2) Normalized Final Distance (NFD) for the distances between objects and the target, measuring partial progress; (3) Waiting Time (WT) for the amount of time an agent waits for assistance with large items; and 4) Action Consistency (AC) for the degree of force alignment when moving items jointly, indicating coordination efficiency. Detailed definitions are in Appx. I.
Human Subject Study  Our study was approved by the IRB. We conducted a human subject study with 32 participants to evaluate BASS against the DP baseline in both tasks. Each participant played 32 maps in total, cooperating with each method in two rounds per map. After completing the first round, the participant and model switched to control the other agent. Upon finishing all maps, participants were given a questionnaire to capture subjective feedback. See Appx. R for details.
Evaluation Protocol Method TCR (↑) NFD (↑) WT (↓) AC (↑) Seen Behaviors MLP 0.2126 0.2987 0.4896 0.8013 GRU 0.2369 0.3011 0.4975 0.8151 MAPPO 0.1929 0.3182 0.5766 0.8097 DP 0.3233 0.5367 0.3789 0.8163 DP/BASS 0.3503 0.5724 0.3598 0.8337 Unseen Behaviors MLP 0.1433 (-32.61%) 0.2413 (-19.22%) 0.5647 (+15.33%) 0.7729 (-3.54%) GRU 0.1638 (-30.87%) 0.2453 (-18.53%) 0.5758 (+15.74%) 0.7830 (-3.94%) MAPPO 0.1635 (-15.19%) 0.2808 (-11.74%) 0.6379 (+10.64%) 0.7858 (-2.95%) DP 0.2563 (-20.72%) 0.4589 (-14.50%) 0.4249 (+12.15%) 0.7854 (-3.78%) DP/BASS 0.3010 (-14.07%) 0.5197 (-9.22%) 0.3899 (+8.37%) 0.8099 (-2.86%) Play with Human DP 0.3855 0.5547 0.4886 0.8054 DP/BASS 0.6512 0.7053 0.3364 0.9124
Table 2: Task 1 results under seen and unseen human behaviors, and with real human partners.
@@ 6.2 Results
Collaboration with Unseen Behaviors (RQ1)  Table 2 reports Task 1 results under three protocols. In the seen setting, agents are trained and evaluated on the same dataset. Here, both DP and BASS outperform MLP, GRU, and MAPPO, with BASS achieving the best task completion (TCR, NFD). In the unseen setting, we split participants into disjoint sets as described in the evaluation protocol, train separate policies, and then evaluate them by playing across groups. All methods degrade when facing unseen behaviors, but BASS shows the least performance drop across TCR, NFD, WT, and AC, indicating stronger robustness. See Appx.S for the full table with standard error.
Collaboration under Unseen Physical Constraints (RQ2)  Fig. 7 shows that, while waiting time and action consistency are comparable across methods, BASS consistently outperforms baselines, particularly in TCR and NFD, which directly reflect task progress under new object properties. This suggests that evaluating candidate actions based on predicted future states helps the model better adapt to variations in size, mass, and shape.
Figure 7: Results on Task 2 under unseen physical constraints.
Collaboration with Humans (RQ3)  Tab. 2 and Fig. 7 show the results with humans. In both tasks, BASS significantly improved task completion rates (TCR and NFD) compared to the DP. This indicates that BASS adapts better to human behavior. For wait time, DP increased when playing with humans, suggesting it struggles with different humans, despite DP capturing multimodal distributions. BASS reduced wait time, demonstrating its ability to adapt to diverse behaviors. For action consistency, DP performed worse because it cannot handle differences between the evaluation and training data. BASS augmented diverse collaborative behaviors during training and selected the best actions for interacting with humans, resulting in better consistency.
@@
Human Feedback (RQ3)  Fig. 8 summarizes post-experiment survey results from humans. We compare BASS with DP. The results show that BASS significantly outperformed DP in the Helpfulness category, indicating that BASS is better at consciously assisting others.
Figure 9: Failure cases: 1) Failing to release items during handover, 2) Not responding when assistance is needed, and 3) Inability to grasp large items upon approach.
Additionally, BASS demonstrated a better understanding of physics, suggesting that our next state predictor effectively comprehends and evaluates different actions to choose the best ones. Independent t-tests revealed that these differences are statistically significant ($p=0.017$).
Failure case study (RQ4)  Fig. 9 shows examples of common failure cases from DP. In task 1, as illustrated in failure case 1, many participants reported that the AI agent frequently holds an item without passing it, resulting in frequent collisions. Additionally, participants noted that the AI agent often failed to come to assist, as shown in failure case 2, where a human agent (blue) was slowly pulling an item, but the AI agent (pink) instead went to grasp other smaller objects. These issues show the model’s limited ability to adapt to diverse behaviors, making it difficult to respond appropriately to actions that were not present in the training dataset. In task 2, most participants pointed out failure case 3, where the AI agent reached the target item but was unable to successfully grasp it. This indicates that the model struggles when encountering objects that were not in the training data. In contrast, BASS shows fewer reported failure cases than DP. Manual inspection revealed that BASS reduced the occurrence rates of the three failure types from {0.797, 0.688, 0.906} in DP to {0.343, 0.563, 0.484}. However, effectively addressing these failures remains a substantial challenge for future research.
@@ 7 Conclusion
We introduce Moving Out, a physically grounded human-AI collaboration benchmark that features a continuous state-action space and dynamic object interactions. We created two challenging tasks and collected human-human collaboration data to enable future model development. Our evaluation results show that much remains to be done with existing models to effectively collaborate with humans in physical environments. Our proposed method, BASS, takes the first step to improve models’ adaptability to diverse human behaviors and physical constraints. Future work includes improving generative models’ ability and inference speed to achieve smooth human-AI collaboration in physical environments, and extending to more complex cooperation dynamics among multiple AI agents and humans.
@@ Appendix A Comparison with Other Environments
Environment State/Action Physics -based Constraints Collaboration Behaviors Overcooked-AI Discrete No Items placed only in specific locations Passing items, dividing tasks, and collision avoidance Watch and Help Discrete Yes Partial observability, diverse objects and goals Goal inference, cooperative help Smart Help Discrete Yes Capability limits (weight, height, open/close/toggle), partial observability Awareness: Goal + capability inference, bottleneck help, avoid unnecessary takeover Table Carrying Continuous No No physical feedback, task ends upon collision Joint carrying (i.e., action consistency) Moving Out Continuous Yes Realistic physics, friction, collision feedback, diverse items with physical properties. Coordination, Awareness of needing help, joint carrying (i.e., action consistency).
Metrics Pros Cons Human Data Number of cooked onions in a limited time Small state/action space, fast training, human data available Limited behavior variety, simple tasks Yes Success Rate, speedup, cumulative reward 3D environment, diverse household tasks No physical variations, high computational cost Synthesized Success Rate (Goal-conditioned), Helping Necessity/Rate, Episode/Success-weighted Length 3D physics, diverse tasks Predefined actions, high computational cost No Success rate, Completion time Continuous actions No physics in interactions, single task, no dataset No Task Completion Rate, Normalized Final Distance, Waiting Time, Action Consistency Realistic physics, multiple collaboration modes, physics feedback, human dataset available Require high-frequency actions for smooth collaboration. Yes
@@
Appendix B Task 1 Data Diversity Analysis
B.1 Visualization of Dataset in Task 1
Figure 10: Visualization of data collected in Task 1 from four groups (each with two players) and from the complete dataset. Blue dots denote the positions of the blue agent and red dots denote the positions of the red agent. The visualizations show clear differences across groups. At the aggregated level, the dataset captures both human behavioral preferences (e.g., preferred paths and object-grasping locations) and broad state coverage.
See Fig. 10.
B.2 Evaluate the diversity of Task 1
To further quantify behavioral diversity in our datasets, we evaluated three distinct sources: the human-human dataset from Moving Out Task 1, a dataset collected from four human experts, and trajectories generated by a trained MAPPO agent. We report diversity across both trajectory- and state-level dimensions. At the trajectory level, we use Dynamic Time Warping (DTW) to compute pairwise distances between all trajectories within a dataset; higher mean and variance indicate greater dissimilarity in path shapes. At the state level, we measure spatial coverage using two complementary metrics: Kernel Density Estimation (KDE) entropy over agent positions and an additional coverage distance metric inspired by [51], which computes the average pairwise distance between trajectories using an RBF kernel. Higher values for both metrics indicate broader exploration of the map. As shown in Table 1, the Task 1 dataset consistently achieves higher scores across all metrics, confirming that data aggregated from 36 human players exhibits substantially greater behavioral diversity compared to expert demonstrations and RL-generated trajectories. This diversity provides a rich foundation for training adaptive collaboration policies and benchmarking generalization.
Appendix C Comparison with Oracle Simulation
@@
C.1 Ablation Study
Task 1 Task 2
Methods TCR↑ NFD↑ TCR↑ NFD↑
GRU 0.3070 0.3674 0.2582 0.3935
+ BASS w/o Simulation 0.4117 0.4396 0.3333 0.4141
+ BASS w/o Augmentation 0.3531 0.4047 0.3670 0.4246
+ Full BASS 0.4120 0.4454 0.3414 0.4410
Diffusion Policy (DP) 0.3829 0.4818 0.3125 0.4526
+ BASS w/o Simulation 0.4028 0.5114 0.3569 0.4908
+ BASS w/o Augmentation 0.4741 0.5561 0.4200 0.5187
+ Full BASS 0.5027 0.5707 0.4348 0.5535
Appendix D Rollout Example
Figure 11 shows an example rollout on Task 2. We only present one example here; additional maps and tasks can be found in the supplementary video.
Figure 11: Comparison of rollouts on Task 2 between our method (BASS), MLP, and DP. The horizontal axis denotes frames as a proxy for time. Our method successfully completes the task at frame 261, whereas both MLP and DP get stuck at an intersection. This highlights the challenge of Task 2, where handling the large fixed-mass circular object is particularly difficult.
Appendix E MAPPO Training Setting
To train MAPPO, we integrate the Moving Out environment into the BenchMARL [52] multi-agent RL library. Our approach to MAPPO training was designed to align with the objectives of Task 1 and Task 2, which were initially conceptualized with dataset-driven methods in mind. We adapted the conditions for MAPPO as follows: